OpenStudy (anonymous):

On a map, Bobby’s house is at (+6, +4) and the pool is at (+3, +4). What is the distance between the points that represent his house and the pool? A. 3 B. 4 C. 6 D. 9

OpenStudy (zepp):

To find the distance between two point, you can use the pythagorean theorem by [ittomg both points in a graph and tracing a triangle, or, you could find it using this formule Distance = \[\sqrt{(x2 - x1)^2+(y2-y1)^2}\] Where x1, y1 is the coordinates of one of the points, and x2, y2 the other one's

OpenStudy (anonymous):

Wat does this even mean?

OpenStudy (zepp):

You plug everything in it, it's just arithmetic x1 = 6 x2 = 3 y1 = 4 y2 = 4 sqrt((3-6)^2+(4-4)^2) = 3

OpenStudy (zepp):

Generally, you find distance between two points with this formula, but this on this one, you should also have used 6-3 = 3, since the y value for both points is the same

OpenStudy (anonymous):

You think its A. 3

OpenStudy (zepp):

Yes.

OpenStudy (anonymous):

Nvm i figured it out in a EASIER way for me but, THANKS!!